OpenStudy (anonymous):

Divide the face of the clock into three parts with two lines so that the sum of the numbers in the three parts are equal.

OpenStudy (anonymous):

I dont get you mathfrick. Shouldnt it be something like this: 1+2+3+...+11+12=78=3*26 So you have to have 26 in each part.

OpenStudy (anonymous):

and you can notice that 12+1=13 and 11+2=13 and so on

OpenStudy (anonymous):

11,12,1,2 9,10,3,4 7,8,5,6

OpenStudy (anonymous):

for getting this draw 2 liines such that u get (11,12,1,2) (9,10,3,4) (5,6,7,8) in each part the sum is 26 and hence all are equal............

OpenStudy (anonymous):

anand is that clear?

OpenStudy (anonymous):

yep that is it!

OpenStudy (anonymous):

yeah!!! u did it...........